OpenStudy (fairy733189):

A building has x floors. The number of apartments on each floor is 6 less than the number of floors. The number of rooms in each apartment is 1 less than the number of apartments on each floor. The expression below shows the total number of rooms in the building: x(x - 6)(x - 7) Part A: What does x(x - 6) represent? Part B: What type of polynomial is x(x - 6)(x - 7)? Part C: What is the degree of the polynomial x(x - 6)(x - 7)?

OpenStudy (fairy733189):

@Mehek14

OpenStudy (ineedhelplz):

x(x-6) represents the no of apartments in building..

OpenStudy (ineedhelplz):

B. x-6=no of aprtments (x-6)-1=x-7=no of rooms in each apartment Third degree

OpenStudy (fairy733189):

Thanks!!

OpenStudy (mertsj):

If you multiply x(x-6)(x-7) you will get: x^3-13x^2+42x The highest exponent is 3 so the degree of the polynomial is 3